# version 0.75

### v4

@@ 2.1.1 Medical Ventilator
In the task of mechanical ventilation, the control algorithm dictates the amount of oxygen inflow $u_{t}$ (in $\mathrm{m}^{3}/s$) into a patient’s lungs. This inflow approximately determines the volume of the lungs $v_{t}$ via a simple linear relation,
@@
where $\Delta$ is the time elapsed between successive actions of the discrete-time control algorithm. Being difficult to directly measure or estimate, the quantity $v_{t}$ is a hidden or latent variable from the controller’s perspective. Instead, the controller observes the air pressure $p_{t}$. We give a simplified physical model relating the pressure and the volume. This model, given below, can be derived by first principles from stress-strain relations, treating the trachea-connected left and right lungs as situated in a two-balloon experiment.
@@
The objective of the control algorithm is to track a physician-prescribed ideal pressure trajectory $p^{*}_{t}$ as closely as possible, say as measured by in squared loss $c_{t}(p_{t},u_{t})=|p_{t}-p^{*}_{t}|^{2}+|u_{t}|^{2}$, while limiting the maximum instantaneous airflow to minimize patient discomfort. See Figure 1.2 for a schematic representation of the airflow when a lung is connected to a mechanical ventilator.
@@ 2.1.2 Double Integrator
The double integrator is a common example of a linear dynamical system. In its simplest form, it models a one-dimensional point object moving according to Newtonian mechanics on a one dimensional trajectory (i.e. the real number line). The state $\mathbf{x}_{t}$ is 2-dimensional, with the first coordinate indicating the position of a point object and the second denoting its velocity. The control algorithm decides the force $u_{t}\in{\mathbb{R}}$ that is applied to the object at each time step. The dynamics equation can be written as
@@
The goal here is to ensure that the point object arrives and then comes to rest at some coordinate $x^{*}\in\mathbb{R}$ as fast as possible, while also not expending too much energy. The cost function is naturally expressed as
@@ 2.1.3 Pendulum Swing-up
The swing-up pendulum is one of the simplest examples of rotational kinematics. It corresponds to a point object of mass $m$ mounted at the end of a massless rod of length $l$. The other end of the rod is fixed to a pivot. Being vertically aligned, the system is subject to gravity with gravitational acceleration $g$. The controller is in charge of applying a time-varying torque $u_{t}\in\mathbb{R}$ at each time step. Given the vertical angle $\theta_{t}$ of the rod, the physics of this system may be summarized as
@@
where $\Delta$ is the time resolution of the controller. Note that this dynamical equation is non-linear in the state $\mathbf{x}_{t}=[\theta_{t},\dot{\theta_{t}}]^{\top}$ due to the presence of the $\sin\theta$ term in the evolution of $\dot{\theta_{t}}$.
Figure 2.3: A pivoted massless rod connected to a point object with mass $m$. The point object is to subject to gravitational acceleration $g$. Source: Wikipedia
The objective here is to get the rod to rest at a vertically upright position at an angle of $\pi$, using as little torque as possible throughout the process. This may be expressed via the cost function
@@ 2.1.4 Aircraft Dynamics
A final example of a dynamical system we consider is the longitudinal (vertical tilt) motion of an aircraft that is regulated by automated controllers. We present the approximate dynamics for the Boeing 747 aircraft, which is a linear dynamical system, from [Boy10]. The 2-dimensional control input $\mathbf{u}_{t}$ corresponds to the elevator angle and the thrust. The disturbances correspond to the velocity of the wind along the body axis and perpendicular to it.
The dynamics equation below has matrices whose numerical values were computed using the dynamics equation for an actual aircraft. These numerical values can be determined by the mass, length, and other physical properties of the aircraft.
@@
Figure 2.4: Longitudinal control in an aircraft. Source: Wikipedia.
The quantities of interest for assessing the performance of the controller are aircraft speed, climb rate, each of which is a linear transformation of the state, and the applied thrust. The corresponding cost is
@@ 2.1.5 Epidemiological Models
The COVID-19 pandemic marked an increased public interest in policy discussions on the effectiveness of certain interventions as a means of controlling the spread of infection. Epidemiological models forecast the spread of an infectious disease via structural models and differential equations. We present the discrete-time analogue of the most basic of such models – the SIR model. Let $S_{t}$, $I_{t}$, and $R_{t}$ denote the susceptible, infected, and recovered population sizes. Let $u_{t}$ be the number of vaccines administered at time $t$. The dynamics of the basic SIR model are given by
@@
Here the coefficient $\beta$ measures how communicable the infection is, $\gamma$ measures how quickly an infected person recovers, and $\alpha$ represents the efficacy of the vaccine in preventing infections.
Figure 2.5: Trajectory of a continuous-time SIR model. Source: Wikipedia.
@@ 2.2 Solution Concepts for Dynamical Systems
Before considering the control, or manipulation, of a dynamical system, we must first reason about what the goal should even be. In this section, we discuss various solution concepts for a dynamical system, including equilibria, stability, stabilizability, and controllability.
Given a dynamical system $\mathbf{x}_{t+1}=f(\mathbf{x}_{t},\mathbf{u}_{t},\mathbf{w}_{t})$, the deterministic-equivalent dynamical system, defined via the dynamics $\mathbf{x}_{t+1}=f^{\prime}(\mathbf{x}_{t},\mathbf{u}_{t})=f(\mathbf{x}_{t},\mathbf{u}_{t},\mathbf{0})$, models the evolution of the dynamical system in the absence of perturbations. Further, the autonomous dynamical system $\mathbf{x}_{t+1}=f^{\prime\prime}(\mathbf{x}_{t})=f(\mathbf{x}_{t},\mathbf{0},\mathbf{0})$ describes state transitions of the dynamical system under the absence of both control inputs and perturbations.
@@ 2.2.1 Equilibrium of Dynamical Systems
An example of an equilibrium is a ball standing at rest at the bottom of a pit. If the ball were to be slightly moved, it would return to the bottom of the pit. Thus, this equilibrium is stable and attractive.
If the ball lies at rest on the top of a mountain, pushing the ball in any direction would cause it to fall. This is an equilibrium which is unstable. A formal definition is as below.
@@
Clearly $\mathbf{0}$ is an equilibrium point as ${A}\mathbf{0}=\mathbf{0}$ for all ${A}$. If the modulus of all eigenvalues is strictly less than one, then after sufficiently many applications of $A$ any vectors converges to the null vector. Thus $\mathbf{0}$ is a globally attractive, stable equilibrium.
In full generality, determining if a system has an equilibrium is an intractable problem, as we show in this chapter.
@@ 2.2.2 Stabilizability and Controllability
A Dynamical System is controllable if and only if, for every initial and target state, there exists a sequence of controls that drives the deterministic-equivalent system to the target state.
Note that controllability clearly implies stabilizability as the zero state is one possible target state.
In the next section, we discuss the computational difficulties that arise in determining if a system is stabilizable and/or controllable.
@@ 2.2.3 Quantitative Definitions of Stabilizability
This definition can apply to any dynamical system and is an important concept in non-stochastic control.
@@ 2.3 Intractability of Equilibrium, Stabilizability, and Controllability
The 3D trajectories of this system are attracted to a 2D manifold with two “wings” that resemble a butterfly as depicted in figure 2.7. For formal properties of this attractor set, see [Str14, Tuc99]. This demonstrates that even studying a solution concept by itself is a formidable task.
@@
Figure 2.6: Julia set. Souce: Wikipedia.
Figure 2.7: Lorenz attractor. Source: Wikipedia.
@@ 2.3.1 Computational Hardness of Stabilizability
Consider the following problem: given a dynamical system, can we determine if it stabilizable, i.e. for every possible starting state does there exist a choice of control inputs that drives the system to the null state? We now show that this decision problem is computationally hard.
@@ 2.4 Bibliographic Remarks
The Lorenz equations were introduced to model the phenomenon of atmospheric convection [Lor63]. For detailed information about the mathematical properties of the Lorenz attractor, see [Str14, Tuc99]. For a rigorous study of Julia sets and the Mandelbrot set, see [BY08].
Nonlinear dynamical systems admit a variety of stability criteria; we chose to define a few that are useful for further developments in the text. A notable omission from our discussion is that of local and global stability; a celebrated result [Lya92] proves that the existence of certain potential functions is sufficient to guarantee global stability, thereby obviating the need for a trajectory-based analysis. An interested reader might look to [Kha15] for an expanded discussion on how these stability criteria compare.
@@ 2.5 Exercises
2.3. Let $\alpha,\beta,\gamma>0$. Does the SIR-model, as it is described in this chapter, define a stabilizable dynamical system (for an appropriate and fixed definition of the “zero state”)? What about controllability? We assume that only starting states with the property $S_{0}+I_{0}+R_{0}=1$ are allowed, and that $S_{t},I_{t},R_{t},u_{t}\geq 0$ has to hold at all times.
@@ 2.4. Consider the double integrator system:
The goal is to ensure the system reaches the desired state $[x^{*},0]^{\top}$ as $t\to\infty$ from any initial condition.
@@
    with $K=[k_{1}\;k_{2}]$, that places all eigenvalues of $A-BK$ strictly inside the unit circle. Find conditions on $\Delta,k_{1},k_{2}$ that ensure all eigenvalues of $A-BK$ are real and lie strictly between $-1$ and $1$, thus guaranteeing $\mathbf{x}_{t}\to[x^{*},0]^{\top}$.
@@
This chapter introduces a setting of equal generality as that of the control problem from the previous chapter: reinforcement learning in Markov Decision Processes (MDP). However, the representation of states and dynamics in a MDP is more explicit and verbose. In such case, we can design efficient methods aimed at optimality given the explicit state and transition representations.
After introducing the reinforcement learning model and Markov Decision Processes, we discuss an important algorithmic technique that will be useful throughout this book: dynamic programming, and the accompanying characterization of optimal policies via the Bellman equation.
@@ 3.1 Reinforcement Learning
Reinforcement learning is a sub-field of machine learning that formally models the setting of learning through interaction in a reactive environment. This is in contrast to learning from examples, as in supervised learning, or inference according to a probabilistic model, as in Bayesian modeling. In reinforcement learning, we have an agent and an environment. The agent observes its position (or state) in the environment and takes actions that transition it to a new state. The environment looks at the agent’s state and hands out rewards based on a set of criteria hidden from the agent.
Figure 3.1: Source: Wikipedia.
Typically, the goal of reinforcement learning is for the agent to learn behavior that maximizes the total reward it receives from the environment. This methodology has led to some notable successes: machines have learned how to play Atari games, and how to beat the highest-rated human players at Go.
@@
The reader can imagine situations where such an assumption is reasonable. For example, in the inverted pendulum setting described in section 2.1.3, the assumption clearly holds. It is not significant for the control problem how a certain state, describing the angle and the angular velocity, was reached, in order to optimally control the pendulum from this state onward.
@@
It can be argued that the Markovian assumption can always be made to work, by adding more features to the state. In the extreme, the state can contain the entire history, which makes the assumption true by default. However, this makes the state representation inefficient.
@@ 3.1.1 A Classical Example: Bubbo’s World
Before diving into the formalism of reinforcement learning, we start with a classical example from the text of Russel and Norvig [RN02]. In this example, a robot starts in the bottom leftmost square, and its goal is to reach the $+1$ reward. The blocks $+1,-1$ represent terminal states, meaning that the game ends when the robot reaches them. The block ‘X’ represents an obstacle through which the robot cannot travel.
@@
In this example, the robot has controls that allow it to move left / right / forward / backward. The controls, however, are not reliable, and with $0.8$ probability the robot moves to wherever it intends. With the remaining $0.2$ probability the robot moves sideways, that is, with probability $0.1$ to each side.
Notice that the Markovian assumption clearly holds for this simple example. What is an optimal solution to this example? A solution concept for a reinforcement learning problem is a mapping from state to action. This mapping is called a “policy”, and it depends on both the structure of the problem as well as on the rewards.
For example, if the rewards at all other states are slightly negative, Bubbo wants to reach the terminal state with a persistent $+1$ reward, as given in the following figure.
@@
We proceed to describe a formal model for capturing this and any other reinforcement learning problem.
@@ 3.2 Markov Decision Processes
In this chapter we describe the general and formal model for reinforcement learning: Markov decision processes.

### v5

@@ 2.1.1 Medical Ventilator
In the task of mechanical ventilation, the control algorithm dictates the amount of oxygen inflow $u_{t}$ (in $\mathrm{m}^{3}/s$) into the lungs of a patient. This inflow approximately determines the volume of the lungs $v_{t}$ via a simple linear relation,
@@
where $\Delta$ is the time elapsed between successive actions of the discrete-time control algorithm. Being difficult to directly measure or estimate, the quantity $v_{t}$ is a hidden or latent variable from the controller’s perspective. Instead, the controller observes the air pressure $p_{t}$. We give a simplified physical model that relates the pressure and the volume. This model, given in the following, can be derived by first principles from stress-strain relations, treating the left and right lungs connected by the trachea as situated in a two-balloon experiment.
@@
The objective of the control algorithm is to follow an ideal pressure trajectory $p^{*}_{t}$ prescribed by the physician as closely as possible, say, as measured by squared loss $c_{t}(p_{t},u_{t})=|p_{t}-p^{*}_{t}|^{2}+|u_{t}|^{2}$, while limiting the maximum instantaneous airflow to minimize patient discomfort. See Figure 1.2 for a schematic representation of airflow when a lung is connected to a mechanical ventilator.
@@ 2.1.2 Double Integrator
The double integrator is a common example of a linear dynamical system. In its simplest form, it models a one-dimensional point object moving according to Newtonian mechanics on a one-dimensional trajectory (i.e. the real number line). The state $\mathbf{x}_{t}$ is 2-dimensional, with the first coordinate indicating the position of a point object and the second denoting its velocity. The control algorithm decides the force $u_{t}\in{\mathbb{R}}$ that is applied to the object at each time step. The dynamics equation can be written as
@@
The goal here is to ensure that the point object arrives and then comes to rest at some coordinate $x^{*}\in\mathbb{R}$ as quickly as possible, while also not expending too much energy. The cost function is naturally expressed as
@@ 2.1.3 Pendulum Swing-up
The swing-up pendulum is one of the simplest examples of rotational kinematics. It corresponds to a point object of mass $m$ mounted at the end of a massless rod of length $l$. The other end of the rod is fixed to a pivot. Being vertically aligned, the system is subject to gravity with gravitational acceleration $g$. The controller is responsible for applying a torque $u_{t}\in\mathbb{R}$ at each time step. Given the vertical angle $\theta_{t}$ of the rod, the physics of this system may be summarized as
@@
where $\Delta$ is the time resolution of the controller. Note that this dynamical equation is non-linear in the state $\mathbf{x}_{t}=[\theta_{t},\dot{\theta_{t}}]^{\top}$ due to the presence of the term $\sin\theta$ in the evolution of $\dot{\theta_{t}}$.
Figure 2.3: A pivoted massless rod connected to a point object with mass $m$. The point object is to subject to gravitational acceleration $g$. Illustration by Krishnavedala, licensed under CC-BY-SA-3.0.
The objective here is to get the rod to rest in a vertical upright position at an angle of $\pi$, using as little torque as possible throughout the process. This may be expressed through the cost function
@@ 2.1.4 Aircraft Dynamics
A final example of a dynamical system we consider is the longitudinal (vertical tilt) motion of an aircraft that is regulated by automated controllers. We present the approximate dynamics for the Boeing 747 aircraft, which is a linear dynamical system, from [Boy10]. The 2-dimensional control input $\mathbf{u}_{t}$ corresponds to the elevation angle and the thrust. The disturbances correspond to the velocity of the wind along the body axis and perpendicular to it.
The following dynamics equation has matrices whose numerical values were computed using the dynamics equation for an actual aircraft. These numerical values can be determined by the mass, length, and other physical properties of the aircraft.
@@
Figure 2.4: Longitudinal control in an aircraft. Graphic by User_A1, licensed under CC-BY-SA 4.0; it is an SVG derivative of a public domain work by G J Vigurs.
The quantities of interest for assessing the performance of the controller are the aircraft speed, climb rate, each of which is a linear transformation of the state, and the applied thrust. The corresponding cost is
@@ 2.1.5 Epidemiological Models
The COVID-19 pandemic marked an increase in public interest in policy discussions on the effectiveness of certain interventions as a means of controlling infection spread. Epidemiological models predict the spread of an infectious disease via structural models and differential equations. We present a discrete-time analogue of the most basic of such models, the SIR model. Let $S_{t}$, $I_{t}$, and $R_{t}$ denote the population sizes of those susceptible, infected, and recovered, respectively. Let $u_{t}$ be the number of vaccines administered at time $t$. The dynamics of the basic SIR model are given by
@@
Here, the coefficient $\beta$ measures the level of communicability of the infection, $\gamma$ measures the rate at which an infected person recovers, and $\alpha$ represents the efficacy of the vaccine in preventing infections.
Figure 2.5: Trajectory of a continuous-time SIR model.
@@ 2.2 Solution Concepts for Dynamical Systems
Before considering the control, or manipulation, of a dynamical system, we must first reason about what the goal should even be. In this section, we discuss various solution concepts for a dynamical system, including equilibrium, stability, stabilizability, and controllability.
Given a dynamical system $\mathbf{x}_{t+1}=f(\mathbf{x}_{t},\mathbf{u}_{t},\mathbf{w}_{t})$, the equivalent deterministic dynamical system, defined through the dynamics $\mathbf{x}_{t+1}=f^{\prime}(\mathbf{x}_{t},\mathbf{u}_{t})=f(\mathbf{x}_{t},\mathbf{u}_{t},\mathbf{0})$, models the evolution of the dynamical system in the absence of perturbations. In addition, the autonomous dynamical system $\mathbf{x}_{t+1}=f^{\prime\prime}(\mathbf{x}_{t})=f(\mathbf{x}_{t},\mathbf{0},\mathbf{0})$ describes the state transitions of the dynamical system in the absence of both control inputs and perturbations.
@@ 2.2.1 Equilibrium of Dynamical Systems
An example of an equilibrium is a ball sitting at rest at the bottom of a pit. If the ball were to be slightly moved, it would return to the bottom of the pit. Thus, this equilibrium is stable and attractive.
If the ball is at rest on the top of a mountain, pushing the ball in any direction would cause it to fall. This is an equilibrium that is unstable. A formal definition is as follows.
@@
Clearly $\mathbf{0}$ is an equilibrium point as ${A}\mathbf{0}=\mathbf{0}$ for all ${A}$. If the modulus of all eigenvalues is strictly less than one, then after sufficiently many applications of $A$ any vector converges to the null vector. Thus, $\mathbf{0}$ is a globally attractive and stable equilibrium.
In general, determining whether a system has an equilibrium is an intractable problem, as we show in this chapter.
@@ 2.2.2 Stabilizability and Controllability
A dynamical system is controllable if and only if, for every initial and target state, there exists a sequence of controls that drives the equivalent deterministic system to the target state.
Note that controllability clearly implies stabilizability, as the zero state is one possible target state.
In the next section, we discuss computational difficulties that arise in determining if a system is stabilizable and/or controllable.
@@ 2.2.3 Quantitative Definitions of Stabilizability
This definition can apply to any dynamical system and is an important concept in nonstochastic control.
@@ 2.3 Intractability of Equilibrium, Stabilizability, and Controllability
The 3D trajectories of this system are attracted to a 2D manifold with two “wings” that resemble a butterfly as depicted in figure 2.7. For formal properties of this attractor set, see [Str14, Tuc99]. This shows that even studying a solution concept itself is a formidable task.
@@
Figure 2.6: Julia set.
Figure 2.7: Lorenz attractor. Visualization by D Schwen, licensed under CC-BY-2.5.
@@ 2.3.1 Computational Hardness of Stabilizability
Consider the following problem: given a dynamical system, can we determine if it stabilizable, i.e., for every possible starting state, does there exist a choice of control inputs that drives the system to the null state? We now show that this decision problem is computationally hard.
@@ 2.4 Bibliographic Remarks
The Lorenz equations were introduced to model the phenomenon of atmospheric convection [Lor63]. For detailed information on the mathematical properties of the Lorenz attractor, see [Str14, Tuc99]. For a rigorous study of Julia sets and the Mandelbrot set, see [BY08].
Nonlinear dynamical systems admit a variety of stability criteria; we chose to define a few that are useful for further developments in the text. A notable omission from our discussion is that of local and global stability; a celebrated result [Lya92] proves that the existence of certain potential functions is sufficient to guarantee global stability, thus obviating the need for a trajectory-based analysis. An interested reader might refer to [Kha15] for an expanded discussion on how these stability criteria compare.
@@ 2.5 Exercises
2.3. Let $\alpha,\beta,\gamma>0$. Does the SIR-model, as is described in this chapter, define a stabilizable dynamical system (for an appropriate and fixed definition of the “zero state”)? What about controllability? We assume that only starting states with the property $S_{0}+I_{0}+R_{0}=1$ are allowed and that $S_{t},I_{t},R_{t},u_{t}\geq 0$ must hold at all times.
@@ 2.4. Consider the double integrator system:
The goal is to ensure that the system reaches the desired state $[x^{*},0]^{\top}$ as $t\to\infty$ from any initial condition.
@@
    with $K=[k_{1}\;k_{2}]$, which places all eigenvalues of $A-BK$ strictly within the unit circle. Find conditions on $\Delta,k_{1},k_{2}$ that ensure that all eigenvalues of $A-BK$ are real and lie strictly between $-1$ and $1$, thus guaranteeing $\mathbf{x}_{t}\to[x^{*},0]^{\top}$.
@@
This chapter introduces a setting of equal generality as that of the control problem from the previous chapter: reinforcement learning in Markov Decision Processes (MDP). However, the representation of states and dynamics in an MDP is more explicit and verbose. In such a case, we can design efficient methods aimed at optimality given the explicit state and transition representations.
After introducing the reinforcement learning model and Markov Decision Processes, we discuss an important algorithmic technique that will be useful throughout this book: dynamic programming and the accompanying characterization of optimal policies via the Bellman equation.
@@ 3.1 Reinforcement Learning
Reinforcement learning is a subfield of machine learning that formally models the setting of learning through interaction in a reactive environment. This is in contrast to learning from examples, as in supervised learning, or inference according to a probabilistic model, as in Bayesian modeling. In reinforcement learning, we have an agent and an environment. The agent observes its position (or state) in the environment and takes actions that transition it to a new state. The environment looks at the agent’s state and hands out rewards based on a set of criteria hidden from the agent.
Figure 3.1: The agent-environment interaction loop in reinforcement learning.
Typically, the goal of reinforcement learning is for the agent to learn behavior that maximizes the total reward it receives from the environment. This methodology has led to some notable successes: machines have learned how to play Atari games and how to beat the highest-rated human players at Go.
@@
The reader can imagine situations where such an assumption is reasonable. For example, in the inverted pendulum setting described in Section 2.1.3, the assumption is clearly valid. It is not significant for the control problem how a certain state, describing the angle and the angular velocity, was reached, in order to optimally control the pendulum from this state onward.
@@
It can be argued that the Markovian assumption can always be made to work by adding more features to the state. In extreme cases, the state can contain the entire history, which makes the assumption true by default. However, this makes the state representation inefficient.
@@ 3.1.1 A Classical Example: Bubbo’s World
Before diving into the formalism of reinforcement learning, we start with a classical example from the text of Russel and Norvig [RN02]. In this example, a robot starts in the bottom left square, and its goal is to reach the $+1$ reward. The blocks $+1,-1$ represent terminal states, which means that the game ends when the robot reaches them. The block ‘X’ represents an obstacle through which the robot cannot travel.
@@
In this example, the robot has controls that allow it to move left / right / forward / backward. However, the controls are not reliable and, with $0.8$ probability, the robot moves wherever it intends to. With the remaining probability $0.2$, the robot moves sideways, that is, with probability $0.1$ to each side.
Notice that the Markovian assumption clearly holds for this simple example. What is an optimal solution to this example? A solution concept for a reinforcement learning problem is a mapping from state to action. This mapping is called a “policy”, and depends both on the structure of the problem and on the rewards.
For example, if the rewards in all other states are slightly negative, Bubbo wants to reach the terminal state with a persistent $+1$ reward, as given in the following figure.
@@
We then describe a formal model to capture this and any other reinforcement learning problem.
@@ 3.2 Markov Decision Processes
In this chapter, we describe the general and formal model for reinforcement learning: Markov decision processes.